\begin{thebibliography}{58}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Ahn et~al.(2016)]{Ahn:2017}
S.~Ahn et~al.
\newblock A neural knowledge language model.
\newblock \emph{arXiv preprint arXiv:1608.00318}, 2016.

\bibitem[Ai et~al.(2016)]{Ai2016doc2veclm}
Q.~Ai et~al.
\newblock Improving language estimation with the paragraph vector model for
  ad-hoc retrieval.
\newblock In \emph{SIGIR}, 2016.

\bibitem[Bansal et~al.(2016)Bansal, Belanger, and A.]{bansal2016ask}
T.~Bansal, D.~Belanger, and M.~A.
\newblock Ask the {GRU}: Multi-task learning for deep text recommendations.
\newblock In \emph{RecSys}, 2016.

\bibitem[Barkan and Koenigstein(2016)]{barkan2016item2vec}
O.~Barkan and N.~Koenigstein.
\newblock {ITEM2VEC}: Neural item embedding for collaborative filtering.
\newblock In \emph{MLSP}, 2016.

\bibitem[Bordes et~al.(2011)Bordes, Weston, Collobert, and
  Bengio]{bordes2011learning}
A.~Bordes, J.~Weston, R.~Collobert, and Y.~Bengio.
\newblock Learning structured embeddings of knowledge bases.
\newblock In \emph{Conference on artificial intelligence}, 2011.

\bibitem[Borisov et~al.(2016)]{Borisov:2016:click}
A.~Borisov et~al.
\newblock A neural click model for web search.
\newblock In \emph{WWW}, 2016.

\bibitem[Cheng et~al.(2016)]{cheng2016wide}
H.~Cheng et~al.
\newblock Wide \& deep learning for recommender systems.
\newblock In \emph{DLRS}, 2016.

\bibitem[Chiu and Nichols(2015)]{Chiu2015ner}
J.~P. Chiu and E.~Nichols.
\newblock Named entity recognition with bidirectional {LSTM}-{CNN}s.
\newblock \emph{TACL}, 2015.

\bibitem[Chuklin et~al.(2015)Chuklin, Markov, and de~Rijke]{chuklin2015click}
A.~Chuklin, I.~Markov, and M.~de~Rijke.
\newblock \emph{Click Models for Web Search}.
\newblock Morgan \& Claypool, 2015.

\bibitem[Collobert and Weston(2008)]{Collobert2008unified}
R.~Collobert and J.~Weston.
\newblock A unified architecture for natural language processing: Deep neural
  networks with multitask learning.
\newblock In \emph{ICML}, 2008.

\bibitem[Dehghani et~al.(2017{\natexlab{a}})Dehghani, Rothe, Alfonseca, and
  Fleury]{Dehghani:CIKM2017}
M.~Dehghani, S.~Rothe, E.~Alfonseca, and P.~Fleury.
\newblock Learning to attend, copy, and generate for session-based query
  suggestion.
\newblock In \emph{CIKM}, 2017{\natexlab{a}}.

\bibitem[Dehghani et~al.(2017{\natexlab{b}})Dehghani, Severyn, Rothe, and
  Kamps]{dehghani2017avoiding}
M.~Dehghani, A.~Severyn, S.~Rothe, and J.~Kamps.
\newblock Avoiding your teacher's mistakes: Training neural networks with
  controlled weak supervision.
\newblock \emph{arXiv preprint arXiv:1711.00313}, 2017{\natexlab{b}}.

\bibitem[Dehghani et~al.(2017{\natexlab{c}})]{Dehghani:2017}
M.~Dehghani et~al.
\newblock Neural ranking models with weak supervision.
\newblock In \emph{SIGIR}, 2017{\natexlab{c}}.

\bibitem[Dietz et~al.(2016)Dietz, Kotov, and Meij]{Dietz2016tutorial}
L.~Dietz, A.~Kotov, and E.~Meij.
\newblock Utilizing knowledge bases in text-centric information retrieval.
\newblock In \emph{ICTIR}, 2016.

\bibitem[Ganguly et~al.(2015)]{Ganguly2015generalizedlm}
D.~Ganguly et~al.
\newblock Word embedding based generalized language model for information
  retrieval.
\newblock In \emph{SIGIR}, 2015.

\bibitem[Graves et~al.(2014)Graves, Wayne, and Danihelka]{Graves:2014:neural}
A.~Graves, G.~Wayne, and I.~Danihelka.
\newblock Neural turing machines.
\newblock \emph{arXiv preprint arXiv:1410.5401}, 2014.

\bibitem[Grbovic et~al.(2015)]{grbovic2015e-commerce}
M.~Grbovic et~al.
\newblock E-commerce in your inbox: Product recommendations at scale.
\newblock In \emph{KDD}, 2015.

\bibitem[Hermann et~al.(2015)]{Hermann:2015:nips}
K.~M. Hermann et~al.
\newblock Teaching machines to read and comprehend.
\newblock In \emph{NIPS}, 2015.

\bibitem[Hewlett et~al.(2016)]{hewlett2016wikireading}
D.~Hewlett et~al.
\newblock {WIKIREADING}: A novel large-scale language understanding task over
  {Wikipedia}.
\newblock In \emph{Proc. ACL}, 2016.

\bibitem[Hidasi et~al.(2016)]{hidashi2016session-based}
B.~Hidasi et~al.
\newblock Session-based recommendations with recurrent neural networks.
\newblock In \emph{ICLR}, 2016.

\bibitem[Huang et~al.(2013)]{huang2013learning}
P.-S. Huang et~al.
\newblock Learning deep structured semantic models for web search using
  clickthrough data.
\newblock In \emph{CIKM}, 2013.

\bibitem[Kannan et~al.(2016)]{kannan2016smart_reply}
A.~Kannan et~al.
\newblock {Smart} {Reply}: Automated response suggestion for email.
\newblock In \emph{KDD}, 2016.

\bibitem[Kenter and de~Rijke(2015)]{kenter2015short}
T.~Kenter and M.~de~Rijke.
\newblock Short text similarity with word embeddings.
\newblock In \emph{CIKM}, 2015.

\bibitem[Kenter and de~Rijke(2017)]{amn2017kenter}
T.~Kenter and M.~de~Rijke.
\newblock Attentive memory networks: Efficient machine reading for
  conversational search.
\newblock In \emph{Workshop on Conversational Approaches to Information
  Retrieval (CAIR'17) at SIGIR 2017}, 2017.

\bibitem[Kenter et~al.(2016)Kenter, Borisov, and
  de~Rijke]{kenter2016siamesecbow}
T.~Kenter, A.~Borisov, and M.~de~Rijke.
\newblock Siamese {CBOW}: Optimizing word embeddings for sentence
  representations.
\newblock In \emph{Proc. ACL}, 2016.

\bibitem[Kenter et~al.(2018)Kenter, Jones, and Hewlett]{kenter2017byte_level}
T.~Kenter, L.~Jones, and D.~Hewlett.
\newblock Byte-level machine reading across morphologically varied languages.
\newblock In \emph{Proc. AAAI}, 2018.

\bibitem[Krizhevsky et~al.(2012)Krizhevsky, Sutskever, and
  Hinton]{krizhevsky2012imagenet}
A.~Krizhevsky, I.~Sutskever, and G.~E. Hinton.
\newblock {ImageNet} classification with deep convolutional neural networks.
\newblock In \emph{NIPS}, 2012.

\bibitem[Lample et~al.(2016)Lample, Ballesteros, Subramanian, Kawakami, and
  Dyer]{Lample2016ner}
G.~Lample, M.~Ballesteros, S.~Subramanian, K.~Kawakami, and C.~Dyer.
\newblock Neural architectures for named entity recognition.
\newblock In \emph{NAACL-HLT}, 2016.

\bibitem[Le and Mikolov(2014)]{le2014distributed}
Q.~V. Le and T.~Mikolov.
\newblock Distributed representations of sentences and documents.
\newblock \emph{arXiv preprint arXiv:1405.4053}, 2014.

\bibitem[Lebret et~al.(2016)Lebret, Grangier, and Auli]{Lebret:2016}
R.~Lebret, D.~Grangier, and M.~Auli.
\newblock Neural text generation from structured data with application to the
  biography domain.
\newblock \emph{arXiv preprint arXiv:1603.07771}, 2016.

\bibitem[Li et~al.(2014)Li, Xu, et~al.]{Li2014semantic}
H.~Li, J.~Xu, et~al.
\newblock Semantic matching in search.
\newblock \emph{FnTIR}, 2014.

\bibitem[Li et~al.(2016)]{li2016deep}
J.~Li et~al.
\newblock Deep reinforcement learning for dialogue generation.
\newblock In \emph{EMNLP}, 2016.

\bibitem[Lu and Li(2013)]{lu2013deep}
Z.~Lu and H.~Li.
\newblock A deep architecture for matching short texts.
\newblock In \emph{NIPS}, 2013.

\bibitem[McAuley et~al.(2015)]{mcauley2015image-based}
J.~McAuley et~al.
\newblock Image-based recommendations on styles and substitutes.
\newblock In \emph{SIGIR}, 2015.

\bibitem[Mei et~al.(2015)Mei, Bansal, and Walter]{mei:2015}
H.~Mei, M.~Bansal, and M.~R. Walter.
\newblock What to talk about and how? {Selective} generation using {LSTM}s with
  coarse-to-fine alignment.
\newblock \emph{arXiv preprint arXiv:1509.00838}, 2015.

\bibitem[Mikolov et~al.(2010)]{mikolov2010recurrent}
T.~Mikolov et~al.
\newblock Recurrent neural network based language model.
\newblock In \emph{Interspeech}, 2010.

\bibitem[Mikolov et~al.(2013)]{mikolov2013efficient}
T.~Mikolov et~al.
\newblock Efficient estimation of word representations in vector space.
\newblock \emph{arXiv preprint arXiv:1301.3781}, 2013.

\bibitem[Mitra et~al.(2017)Mitra, Diaz, and Craswell]{Mitra:2016:www}
B.~Mitra, F.~Diaz, and N.~Craswell.
\newblock Learning to match using local and distributed representations of text
  for web search.
\newblock In \emph{WWW '17}, 2017.

\bibitem[Mitra et~al.(2016)]{mitra2016desm}
B.~Mitra et~al.
\newblock A dual embedding space model for document ranking.
\newblock \emph{arXiv preprint arXiv:1602.01137}, 2016.

\bibitem[Oord et~al.(2013)Oord, Dieleman, and Schrauwen]{oord2013deep}
A.~Oord, S.~Dieleman, and B.~Schrauwen.
\newblock Deep content-based music recommendation.
\newblock In \emph{NIPS}, 2013.

\bibitem[Quadrana et~al.(2017)]{quadrana2017personalizing}
M.~Quadrana et~al.
\newblock Personalizing session-based recommendations with hierarchical
  recurrent neural networks.
\newblock In \emph{RecSys}, 2017.

\bibitem[Robertson et~al.(1995)]{robertson1995okapi}
S.~E. Robertson et~al.
\newblock Okapi at {TREC-3}.
\newblock \emph{Nist Special Publication Sp}, 1995.

\bibitem[Salakhutdinov and Hinton(2009)]{salakhutdinov2009semantic}
R.~Salakhutdinov and G.~Hinton.
\newblock Semantic hashing.
\newblock \emph{International Journal of Approximate Reasoning}, 2009.

\bibitem[Serban et~al.(2016)]{serban2016generating}
I.~V. Serban et~al.
\newblock Generating factoid questions with recurrent neural networks: The 30m
  factoid question-answer corpus.
\newblock \emph{arXiv preprint arXiv:1603.06807}, 2016.

\bibitem[Sutskever et~al.(2014)Sutskever, Vinyals, and
  Le]{sutskever2014sequence}
I.~Sutskever, O.~Vinyals, and Q.~V. Le.
\newblock Sequence to sequence learning with neural networks.
\newblock In \emph{NIPS}, 2014.

\bibitem[Szummer and Yilmaz(2011)]{szummer2011semi}
M.~Szummer and E.~Yilmaz.
\newblock Semi-supervised learning to rank with preference regularization.
\newblock In \emph{CIKM}, 2011.

\bibitem[Van~Gysel et~al.(2016{\natexlab{a}})Van~Gysel, de~Rijke, and
  Kanoulas]{VanGysel:2016:cikm}
C.~Van~Gysel, M.~de~Rijke, and E.~Kanoulas.
\newblock Learning latent vector spaces for product search.
\newblock In \emph{CIKM}, 2016{\natexlab{a}}.

\bibitem[Van~Gysel et~al.(2016{\natexlab{b}})Van~Gysel, de~Rijke, and
  Worring]{VanGysel:2016:www}
C.~Van~Gysel, M.~de~Rijke, and M.~Worring.
\newblock Unsupervised, efficient and semantic expertise retrieval.
\newblock In \emph{WWW}, 2016{\natexlab{b}}.

\bibitem[Van~Gysel et~al.(2017{\natexlab{a}})Van~Gysel, de~Rijke, and
  Kanoulas]{VanGysel2017nvsm}
C.~Van~Gysel, M.~de~Rijke, and E.~Kanoulas.
\newblock Neural vector spaces for unsupervised information retrieval.
\newblock \emph{arXiv preprint arXiv:1708.02702}, 2017{\natexlab{a}}.

\bibitem[Van~Gysel et~al.(2017{\natexlab{b}})]{VanGysel2017proactive}
C.~Van~Gysel et~al.
\newblock Reply with: Proactive recommendation of email attachments.
\newblock In \emph{CIKM}, 2017{\natexlab{b}}.

\bibitem[Vasile et~al.(2016)]{vasile2016meta-prod2vec}
F.~Vasile et~al.
\newblock {Meta-Prod2Vec} -- product embeddings using side-information for
  recommendations.
\newblock In \emph{RecSys}, 2016.

\bibitem[Vinyals and Le(2015)]{vinyals2015neural}
O.~Vinyals and Q.~Le.
\newblock A neural conversational model.
\newblock In \emph{ICML}, 2015.

\bibitem[Vuli{\'c} and Moens(2015)]{vulic2015monolingual}
I.~Vuli{\'c} and M.-F. Moens.
\newblock Monolingual and cross-lingual information retrieval models based on
  (bilingual) word embeddings.
\newblock In \emph{Proc. SIGIR}, 2015.

\bibitem[Wang et~al.(2015)]{wang2015collaborative}
H.~Wang et~al.
\newblock Collaborative deep learning for recommender systems.
\newblock In \emph{KDD}, 2015.

\bibitem[Wang et~al.(2014)Wang, Zhang, Feng, and Chen]{Wang2014knowledge}
Z.~Wang, J.~Zhang, J.~Feng, and Z.~Chen.
\newblock Knowledge graph embedding by translating on hyperplanes.
\newblock In \emph{Proc. AAAI}, 2014.

\bibitem[Weston et~al.(2016)]{weston2015towards}
J.~Weston et~al.
\newblock Towards {AI}-complete question answering: A set of prerequisite toy
  tasks.
\newblock In \emph{ICLR}, 2016.

\bibitem[Zhao et~al.(2015)Zhao, Zhiyuan, and Sun]{Zhao2015}
Y.~Zhao, L.~Zhiyuan, and M.~Sun.
\newblock Representation learning for measuring entity relatedness with rich
  information.
\newblock In \emph{IJCAI}, 2015.

\bibitem[Zuccon et~al.(2015)Zuccon, Koopman, Bruza, and
  Azzopardi]{Zuccon2015nntm}
G.~Zuccon, B.~Koopman, P.~Bruza, and L.~Azzopardi.
\newblock Integrating and evaluating neural word embeddings in information
  retrieval.
\newblock In \emph{ADCS}, 2015.

\end{thebibliography}

\documentclass[sigconf]{acmart}

\usepackage{booktabs} 
\usepackage[skip=0pt]{caption} 

\setcopyright{rightsretained}

\acmConference[WSDM 2018]{WSDM 2018}{February 5--9, 2018}{Los Angeles, USA} 
\acmYear{2018}
\copyrightyear{2018}

\acmPrice{15.00}

\usepackage[subtle, wordspacing=normal, tracking=normal, bibnotes, charwidths=normal, leading, indent, lists, paragraphs=normal, mathspacing=normal, bibliography=normal]{savetrees}
\usepackage{flushend}

\usepackage{multirow}
\usepackage{enumitem}

\begin{document}

\title{Neural Networks for Information Retrieval}
\titlenote{Slides and other materials will be made available online at \url{http://nn4ir.com}.}

\author{Tom Kenter}
\authornote{Corresponding author.}
\affiliation{%
  \institution{Booking.com}
  \city{Amsterdam, The Netherlands}
}
\email{tom.kenter@gmail.com}

\author{Alexey Borisov}
\affiliation{%
  \institution{Yandex}
  \city{Moscow, Russia}
}
\email{alborisov@yandex-team.ru}

\author{Christophe Van Gysel}
\affiliation{%
  \institution{University of Amsterdam}
  \city{Amsterdam, The Netherlands}
}
\email{cvangysel@uva.nl}

\author{Mostafa Dehghani}
\affiliation{%
  \institution{University of Amsterdam}
  \city{Amsterdam, The Netherlands}
}
\email{dehghani@uva.nl}

\author{Maarten de Rijke}
\affiliation{%
  \institution{University of Amsterdam}
  \city{Amsterdam, The Netherlands}
}
\email{derijke@uva.nl}

\author{Bhaskar Mitra}
\affiliation{%
  \institution{Microsoft, University College London}
  \city{Cambridge, UK}
}
\email{bmitra@microsoft.com}

\renewcommand{\shortauthors}{T. Kenter et al.}

\settopmatter{printacmref=true}
\fancyhead{}

\copyrightyear{2018} 
\acmYear{2018} 
\setcopyright{rightsretained} 
\acmConference[WSDM 2018]{WSDM 2018: The Eleventh ACM International Conference on Web Search and Data Mining }{February 5--9, 2018}{Marina Del Rey, CA, USA}
\acmBooktitle{WSDM 2018: WSDM 2018: The Eleventh ACM International Conference on Web Search and Data Mining , February 5--9, 2018, Marina Del Rey, CA, USA}
\acmDOI{10.1145/3159652.3162009}
\acmISBN{978-1-4503-5581-0/18/02}

\maketitle


\subsection*{Abstract}
Machine learning plays a role in many aspects of modern IR systems, and deep learning is applied in all of them.
The fast pace of modern-day research has given rise to many approaches to many IR problems.
The amount of information available can be overwhelming both for junior students and for experienced researchers looking for new research topics and directions.
The aim of this full-day tutorial is to give a clear overview of current tried-and-trusted neural methods in IR and how they benefit IR.


\subsection*{Motivation}

Prompted by advances of deep learning in computer vision, neural networks (NNs) have resurfaced as a popular machine learning paradigm in many other directions of research, including IR.
Recent years have seen NNs being applied to all key parts of the typical modern IR pipeline, such as 
click models, 
core ranking algorithms, 
dialogue systems,
entity retrieval, 
knowledge graphs, 
language modeling, 
question answering, 
and
text similarity. 

An advantage that sets NNs apart from many learning strategies employed earlier, is their ability to work from raw input data.
Where designing features used to be a crucial aspect of newly proposed IR approaches, the focus has shifted to designing network architectures instead.
Hence, many architectures and paradigms have been proposed, such as auto-encoders, recursive networks, recurrent networks, convolutional networks, various embedding methods, and deep reinforcement learning.
This tutorial aims to provide an overview of the main network architectures currently applied in IR and to show how they relate to previous work.

We expect the tutorial to be useful both for academic and industrial researchers and practitioners who either want to develop new neural models, use them in their own research in other areas or apply the models described here to improve actual IR systems.


\subsection*{Brief outline of the topics to be covered}
Table~\ref{table:schedule} gives an overview of the time schedule of the tutorial.
The total time is 6 hours, plus breaks.
\begin{table}[h]
  \centering
  \caption{Time schedule for NN4IR tutorial}
  \label{table:schedule}
  \begin{tabular}{@{}l l l l@{}}
    \toprule
    \multicolumn{2}{c}{Morning} & \multicolumn{2}{c}{Afternoon} \\
    \cmidrule(r){1-2}\cmidrule(l){3-4}
    Preliminaries          & 45 min. &  Recommender systems    & 45 min. \\
    Semantic matching      & 45 min. &     Modeling user behavior & 45 min. \\
    Learning to rank       & 45 min. &     Generating responses   & 45 min. \\
    Entities               & 45 min.  &   Industry insights      & 45 min. \\
   \bottomrule
  \end{tabular}
\end{table}
We bring a team of six lecturers, all with their specific areas of specialization.
Each session will have two expert lecturers (indicated by their initials below) who will together present the session.

{
\setlength{\baselineskip}{3pt}


\paragraph{\textbf{Preliminaries} [TK, MdR]}
The recent surge of interest in deep learning has given rise to a myriad of model architectures.
Different though the inner structures of NNs can be, many building blocks are shared.
In this preliminary session, we focus on key concepts, all of which will be referred to multiple times in subsequent sessions.
In particular we will cover distributed representations/embeddings \cite{mikolov2013efficient}, fully-connect layers, convolutional layers \cite{krizhevsky2012imagenet}, recurrent networks \cite{mikolov2010recurrent} and sequence-to-sequence models \cite{sutskever2014sequence}.


\paragraph{\textbf{Semantic matching} [CVG, BM]}
The problem of matching based on textual descriptions arises in many retrieval systems. The traditional IR approach involves computing lexical term overlap between query and document \citep{robertson1995okapi}. However, a vocabulary gap occurs when query and documents use different terms to describe the same concepts \citep{Li2014semantic}.
Semantic matching methods bridge the vocabulary gap by matching concepts rather than exact word occurrences.
We cover neural network-based methods that learn to provide a semantic matching signal supervised fashion \citep{lu2013deep,Mitra:2016:www,huang2013learning,mitra2016desm}, semi-supervised \citep{Dehghani:2017, dehghani2017avoiding}, and unsupervised \citep{vulic2015monolingual,kenter2015short,Zuccon2015nntm,Ganguly2015generalizedlm,le2014distributed,VanGysel2017nvsm,VanGysel:2016:cikm,VanGysel:2016:www,Ai2016doc2veclm,kenter2016siamesecbow}.


\paragraph{\textbf{Learning to rank} [AB, MD]}
Capturing the notion of relevance for ranking needs to account for different aspects of the query, the document, and their relationship.
Neural methods for ranking can use manually crafted query and document features, and combine them with regards to a ranking objective.
Moreover, latent representations of the query and document can be learned in situ. 
We cover scenarios with different levels of supervision---unsuper\-vised \cite{VanGysel:2016:www, VanGysel:2016:cikm, salakhutdinov2009semantic}, semi/weakly-supervised \cite{Dehghani:2017,dehghani2017avoiding, szummer2011semi}, or fully-supervised using labeled data \cite{Mitra:2016:www} or interaction data \cite{huang2013learning}.


\paragraph{\textbf{Entities} [CVG, TK]}
Entities play a central role in modern IR systems \citep{Dietz2016tutorial}.
We cover neural approaches to solving the basic task of named entity recognition~\citep{Collobert2008unified,Chiu2015ner,Lample2016ner}, as well learning representations in an end-to-end neural model for learning a specific task like entity ranking for expert finding~\citep{VanGysel:2016:www}, product search~\citep{VanGysel:2016:cikm} or email attachment retrieval \citep{VanGysel2017proactive}.
Furthermore, work related to knowledge graphs will be covered, such as graph embeddings \citep{bordes2011learning,Wang2014knowledge,Zhao2015}.


\paragraph{\textbf{Recommender systems} [MdR, BM]}
Deep learning has also found its way into recommender systems.
We cover learning of item (products, users) embeddings~\citep{barkan2016item2vec,grbovic2015e-commerce,vasile2016meta-prod2vec}, as well as deep collaborative filtering using different deep learning techniques and architectures~\citep{wang2015collaborative,cheng2016wide}.
Furthermore, NN-based feature extraction from content (such as images, music, text)~\citep{bansal2016ask,oord2013deep,mcauley2015image-based}, and session-based recommendations with RNNs~\citep{Dehghani:CIKM2017, hidashi2016session-based,quadrana2017personalizing} will be covered.

\paragraph{\textbf{Modeling user behavior} [AB, MdR]}
Modeling user browsing behavior plays an important role in modern IR systems.
Accurately interpreting user clicks is difficult due to various types of bias.
Many click models based on Probabilistic Graphical Models (PGMs) have been proposed~\cite{chuklin2015click}.
Recently, it was shown that recurrent neural networks can learn to account for biases in user clicks directly from click-through, i.e., without the need for a predefined set of rules as is customary for PGM-based click models~\cite{Borisov:2016:click}.


\paragraph{\textbf{Generating responses} [TK, MD]}

Recent inventions such as smart home devices, voice search, and virtual assistants provide new ways of accessing information.
They require a different response format than the classic ten blue links.
Examples are conversational and dialog systems \citep{li2016deep,vinyals2015neural} or machine reading and question answering tasks, where the model either infers the answer from unstructured text \citep{Hermann:2015:nips, weston2015towards,hewlett2016wikireading,amn2017kenter,serban2016generating,kenter2017byte_level}  or generates natural language given structured data, like data from knowledge graphs or from external memories~\citep{Graves:2014:neural, Ahn:2017,Lebret:2016, mei:2015}.


\paragraph{\textbf{Industry insights} [AB, BM]}
Where the focus of academic papers can be on a specific subtask, industry approaches have to ensure that a system works from start to end.
As a result, extra challenges are involved concerning the user experience.
For example in Google's SmartReply system \cite{kannan2016smart_reply} the neural model at the core of the system is embedded in a much larger framework of non-neural methods to make sure quality and efficiency requirements are met.

}

\bibliographystyle{abbrvnatnourl}
\bibliography{wsdm-tutorial-nn4ir} 

\end{document}